# The self-interaction effects on the Kerr black hole superradiance and their observational implications

### v1

@@ VI CONCLUSION
Based on these results, we discuss the impact of self-interactions on gravitational waves emitted from boson cloud annihilation in gravitational atoms, including both frequency and amplitude modifications. The predicted frequency drift of the emitted gravitational waves could be detectable by current or future observatories. Moreover, our calculations suggest that, for gravitational atoms composed of scalar boson clouds, the black hole may retain a significant fraction of its initial mass and spin even after the dominant superradiant mode terminates if self-interactions are present. This finding could substantially relax current constraints on scalar boson models derived from black hole spin measurements. As specific examples, we use GW190412 and GW190517 to demonstrate how self-interaction can alleviate these constraints, providing quantitative estimates of the self-interaction energy scale required to do so. Additionally, our results can help explain several existing experimental measurements of black hole spins.
Overall, the conclusions drawn from this work provide valuable insights for future dark matter searches using gravitational wave observations and for astrophysical constraints on dark matter properties based on black hole observations. In investigating the superradiance of self-interacting scalar bosons around rotating black holes, we solve the Klein-Gordon equation in the Kerr metric and, for the first time in a relativistic framework, apply the continued fraction method to compute the eigenfrequencies of boson clouds with self-interactions in gravitational atoms. This approach allows us to determine the superradiant growth rates of self-interacting bosons for specific energy levels around Kerr black holes and, in principle, to numerically solve for the growth rates at any stage of the boson cloud formation across the entire parameter space of $\alpha$ and $f_{a}$. In the limit where the self-interaction scale $f_{a}$ approaches infinity, our results recover those of the non-interacting case.
Nevertheless, our calculations assume that the boson cloud is dominated by a single energy level, neglecting potential level mixing between different states and the influence of higher bound states on the dominant mode. Future work could expand on this study in several directions. For example, one could consider higher-order self-interaction terms beyond the quartic term, incorporating their effects within the current methodological framework. Additionally, the impact of self-interactions on the annihilation of boson pairs into gravitons within the cloud could be explored in greater detail, providing a more comprehensive understanding of gravitational wave signals from gravitational atoms. From a methodological perspective, applying numerical relativity to dynamically model the formation and evolution of self-interacting boson clouds around black holes could yield more precise results for system parameters such as growth rates and offer validation for the methods employed in this study.
@@ Acknowledgements.
We thank Junquan Su, Siyu Jiang, Jing Yang and Xiang’en Yang for their useful suggestions on the manuscript. This work was supported by the National Natural Science Foundation of China (NNSFC) under Grant No. 12205387 and No. 12475111.

### v2

@@ VI Conclusions and discussions
Based on these results, we discuss the impact of self-interactions on GWs emitted from boson cloud annihilation in gravitational atoms, including both frequency and amplitude modifications. The predicted frequency drift of the emitted GWs could be detectable by current or future observatories. Moreover, our calculations suggest that, for gravitational atoms composed of scalar boson clouds, the BH may retain a significant fraction of its initial mass and spin even after the dominant superradiant mode terminates if self-interactions are present. This finding could substantially relax current constraints on scalar boson models derived from BH spin measurements. As specific examples, we use GW190412 and GW190517 to demonstrate how self-interaction can alleviate these constraints, providing quantitative estimates of the self-interaction energy scale required to do so. Additionally, our results can help explain several existing experimental measurements of BH spins.
Overall, the conclusions drawn from this work provide valuable insights for future dark matter searches using GW observations and for astrophysical constraints on dark matter properties based on BH observations. In investigating the superradiance of self-interacting scalar bosons around rotating BHs, we solve the Klein-Gordon equation on the Kerr metric and, for the first time in a relativistic framework, apply the continued fraction method to compute the eigenfrequencies of boson clouds with self-interactions in gravitational atoms. This approach allows us to determine the superradiant growth rates of self-interacting bosons for specific energy levels around Kerr BHs and, in principle, to numerically solve for the growth rates at any stage of the boson cloud formation across the entire parameter space of $\alpha$ and $f_{a}$. In the limit where the self-interaction scale $f_{a}$ approaches infinity, our results recover those of the non-interacting case.
Nevertheless, our calculations assume that the boson cloud is dominated by a single energy level, neglecting mode mixing between different states and the influence of higher bound states on the dominant mode. Future work could expand upon this study in several directions. For example, one could consider higher-order self-interaction terms beyond the quartic term, incorporating their effects within the current methodological framework. Additionally, the impact of self-interactions on the annihilation of boson pairs into gravitons within the cloud could be explored in greater detail, providing a more comprehensive understanding of GW signals from gravitational atoms. From a methodological perspective, applying numerical relativity to dynamically model the formation and evolution of self-interacting boson clouds around a BH could yield more precise results for system parameters such as growth rates and validate the methods employed in this study.
@@ Acknowledgements.
We thank Siyu Jiang, Dayun Qiu, Jing Yang, Xiang’en Yang, and Junquan Su for their useful suggestions on the manuscript. This work was supported by the National Natural Science Foundation of China (NNSFC) under Grants No. 12475111 and No. 12205387.